OpenStudy (anonymous):

Find the indicated value. Find f(4) when f(x)= 5x^2+4x a) 64 b) 36 c)96 d)56

OpenStudy (paxpolaris):

When x=4, \[5x^2+4x = 5\left( 4^2 \right)+4 \cdot 4\]

OpenStudy (paxpolaris):

\[= 5\cdot 16 + 16\]

OpenStudy (anonymous):

So C 96